A map has the scale 1 inch = 10 miles. On the map, the area of Big Bend National Park in Texas is about 12.5 square inches. Estimate the actual area of the park in acres. (Hint: 1 square mile = 640 acres)
Answer by jim_thompson5910(35256)

1 inch = 10 miles

(1 inch)^2 = (10 miles)^2

1 square inch = 100 square miles

-------------------------------------------------

Convert 12.5 square inches on the map to square miles in real life

(12.5 square inches) * [ (100 square miles)/(1 square inch) ] = 12.5*100 = 1250

The park is approximately 1250 square miles

-------------------------------------------------

Now convert to acres

(1250 square miles) * [ (640 acres)/(1 square mile) ] = 1250*640 = 800,000

The park is approximately 800,000 acres